OpenStudy (anonymous):

larry is trying to figure out the weights of 3 objects. here are all the facts he knows. -the sum of all the objects is 28 pounds -the heaviest one is 15 pounds 2 ounces -the other 2 objects differ in weight by 4 ounces. Please help!

OpenStudy (anonymous):

DAWWG. I TOTES DEF HAD THIS PROBLEME TOOO.

OpenStudy (anonymous):

Assign variables to each object, then create equations out of what you know.

OpenStudy (anonymous):

x +y + z = 28 pounds z = 15 pounds 2 ounces (I don't know imperical so you'll have to convert ounces into decimals) y - x = 4 ounces (which is 0.something pounds, again, I just work in metric)

OpenStudy (anonymous):

wait so whats the answer??

OpenStudy (anonymous):

I don't know, but that's how you get to it.

OpenStudy (anonymous):

If y - x = 4 ounces, then you can also say 4 + x = y... so go back to the first equation and put that in for y. So, now you have x + (4 + x) +z = 28 pounds

OpenStudy (anonymous):

You know the value for z, so plug that in and you have x + (4 ounces + x) + 15 pounds 2 ounces = 28 pounds. Then, if you simpify, you have 2x + 15 pounds 6 ounces = 28 pounds. Solve for x, then you can solve for y, and then you're done!

OpenStudy (dumbcow):

16 ounces in a pound z = 15.125 y-x = 0.25

OpenStudy (anonymous):

Thank god, someone who knows imperical! *phew*

OpenStudy (anonymous):

i still have no idea what the answer is?

OpenStudy (anonymous):

Okay, we know z = 15.125, y - x = 0.25, and x + y + z = 28, right?

OpenStudy (anonymous):

You can substitute z in the third equation, and it becomes x + y + 15.125 = 28. What we want to do is to try to get rid of all the variables. So, you take the second equation I mentioned and move it around until you can solve for y. That would be 0.25 + x = y.

OpenStudy (anonymous):

So, now you know y (sort of) and you, again, plug it into the last equation. Instead of x + y + 15.125 = 28, we've got x + (0.25 + x) + 15.125 = 28. Simplified, it's 2x + 15.375 = 28. Solve for x, and then you know the weights of all of the objects.

OpenStudy (anonymous):

x = the weight of the lightest object, y = the weight of the middle object, z = the weight of the heaviest object.